OpenStudy (anonymous):

solve for y in the first equation. x + 3y = 6 2x – y = 10

OpenStudy (anonymous):

I hope I'm doing this right :o) x + 3y = 6---> Multiply by -2 2x – y = 10 -2(x+3y=6)--> -2x-6y=-12 -2x-6y=-12 +2x- y = 10 ------------ 0x -5y = -2 -5y= -2 /-5 /-5 y=0.4 x+3(0.4)=6 x+1.2=6 x=4.8 4.8+3(0.4)=6 4.8+1.2=6 6=6 ✔